# Highfield, Southampton S017 1BJ, United Kingdom

### v1

@@ 1 Introduction
[…]
Beyond the original motivation of the axion solution to the strong CP problem Peccei:1977hh ; Peccei:1977ur ; Wilczek:1977pj ; PhysRevLett.40.279 , ALPs can arise in many BSM scenarios such as string compactifications Cicoli:2013ana ; Ringwald:2012cu , supersymmetric theories  Bellazzini:2017neg , neutrino-mass-generation mechanisms Alexandre:2020tba ; Mavromatos:2020hfy , as well as composite-Higgs realisations featuring additional pNGBs Gripaios:2009pe . In addition, ALPs can be viable dark matter candidates as well as mediators to the dark-matter sector, see e.g. Ref. Fitzpatrick:2023xks ; Dror:2023fyd .
@@
Without committing to a specific UV completion, we will study the phenomenology of the top-philic ALP at the LHC, focusing on the model-independent features that primarily depend on the low-energy physics described by the EFT (see Ebadi:2019gij ; Esser:2023fdo ; Rygaard:2023dlx ; Bruggisser:2023npd for recent related studies). The first observation is that the top-philic nature of the ALP is not preserved beyond tree level. In fact $c_{t}$ induces, through radiative corrections involving SM fields, couplings to all SM fermions and gauge bosons. As we will explore and discuss in detail, this has important implications for the possible discovery channels of this light BSM state. For example, for ALP masses below the bottom threshold, the constraints on the model are dominated by flavour observables (see, e.g., Ref. Bauer:2021mvw ) from the induced fermion couplings. At even lighter, keV-scale masses, the induced coupling to electrons implies extremely strong bounds from stellar cooling via ALP emission, beyond any conceivable collider sensitivity Chala:2020wvs . Above the top-quark pair threshold, a top-philic ALP can be probed efficiently by resonant $t\bar{t}$ searches. However, by this point the ALP can no longer be considered light and shift-symmetry-breaking effects start to become significant.
@@ 2.1 The ALP effective Lagrangian
$\psi_{f}\rightarrow{\rm exp}\left(ic\frac{a(x)}{f_{a}}Q_{f}\right)\psi_{f},\quad\phi\rightarrow{\rm exp}\left(ic\frac{a(x)}{f_{a}}Q_{H}\right)\phi.$ (3)
@@ 2.2.2 Coupling to gluons and photons
The large suppression in Eq. (13) for relatively light ALPs raises the question of whether higher-loop contributions involving lighter fermions can actually become the dominant ones for $m_{a}\ll m_{t}$. While a full two-loop computation of $\mathcal{A}^{\mu\nu}(a\rightarrow gg)$ is beyond the scope of this work, we can estimate the size of this contribution. To this end, we consider the one-loop diagrams involving light fermions that couple to the ALP with the effective (one-loop) couplings in Eq. (9) and Eq. (10). These contributions correspond to diagrams such as those in Fig. 2. However, one should keep in mind that at the same perturbative order also diagrams like that represented in Fig. 2, which we do not take into account, are present.
Including the extra contributions from diagrams like the one in Fig. 2, the effective interaction of Eq. (13) generalises to
@@
For the range we will consider in our work, $10\leavevmode\nobreak\ {\rm GeV}\lesssim m_{a}\lesssim 200\leavevmode\nobreak\ {\rm GeV}$, Eq. (17) shows that contributions from up, down and strange quarks are negligible, while the charm mass starts to be relevant for $m_{a}\sim 10\leavevmode\nobreak\ {\rm GeV}$. Also, in the range $2m_{b}\ll m_{a}\ll 2m_{t}$, combining Eqs. (9), (16) and (17) one obtains
@@
First we notice, that at small $m_{a}$ values the bottom-quark contribution, with $c_{b}$ obtained from one-loop induced $c_{t}$ effects, is dominant w.r.t. the top-quark one. For larger $m_{a}$, the top contribution takes over. Second, in the mass-window of phenomenological interest in our paper, $10\leavevmode\nobreak\ {\rm GeV}\lesssim m_{a}\lesssim 200\leavevmode\nobreak\ {\rm GeV}$, $c_{GG}^{\rm eff}$ never surpasses 0.05, contrary to the $O(1)$ expectations from typical ALP models. Finally, there is a cancellation between the top and bottom contributions, arising at one-and two-loop level, respectively. We will discuss this in Sec. 3.1 and in Appendix E, noting that it does not impact the exclusion limits that will be derived in this work.
@@ 3.1 Branching ratios
The decay into photons follows a pattern very similar to the one discussed for the case of the decay into gluons. However, in this case, the loop-induced contributions involve electric charges and hence the minimum of the BR occurs at a numerically different value of $m_{a}$ w.r.t. the decay into gluons. The same argument regarding the location of the minimum and the irrelevance for our analysis applies to this decay channel. In particular, the relevant information for our analysis is that the BR of this channel is always (much) smaller than 0.1%. We also observe that, unlike the case of the gluons, there is no small change increase in the region close to $m_{a}\approx 2m_{b}$; rather, there is an even stronger decrease, due to the additional dependence on the electric charge of the light fermion in the loop.
@@ 3.2 LHC production cross sections
Process
Sub-processes $(Q=c,b)$
Scale Cuts
$pp\rightarrow a$
$gg\rightarrow a$, $Q\bar{Q}\rightarrow a$
$m_{a}$ None
$pp\rightarrow a+j$
$gg\rightarrow ga$, $gq\rightarrow qa$, $q\bar{q}\rightarrow ga$, $Q\bar{Q}\rightarrow ga$
$m_{\rm T}(a)$ $p_{\rm T}(a)>30\,{\rm GeV},\,|\eta(a)|<4$
$pp\rightarrow a+j_{h}$
$gQ\rightarrow Qa$
$m_{\rm T}(a)$ $p_{\rm T}(a)>30\,{\rm GeV},\,|\eta(a)|<2.5$
$pp\rightarrow t\bar{t}a$
$gg\rightarrow t\bar{t}a$, $q\bar{q}\rightarrow t\bar{t}a$
$m_{t}+\frac{m_{a}}{2}$ None
@@
As in the case of $gg\rightarrow a$, in addition to the contributions induced by $c_{t}$, the contributions from $c_{f}$ associated to $f\neq t$ cannot be neglected. These correspond to an approximate two-loop calculation that we obtain by plugging in the $c_{f}$ generated at one-loop. Concerning the change of basis from derivative to non-derivative, it is crucial to ensure that the definition of $\tilde{c}_{GG}$ in Eq. (35) includes $c_{f}$ for $f\neq t$ consistently with the $f$ running in the loop diagrams such as as the one on the left of Fig. 5. Unlike the case of direct ALP production or its decay into gluons or photons, however, predictions for the $pp\rightarrow a+j$ production are much more reliable. Indeed, at high $p_{T}(a)$, which is the phenomenologically relevant kinematic configuration in our work, one expects a contribution from the top-quark loops of the form given in Eqs. (19) and (20) and a contribution from the other fermions of the form given in (21). Notice that, in the non-derivative basis, this means that the purely pseudoscalar component, associated to the left-hand diagram in Fig. 5 vanishes and all of the contribution is given by the tree-level diagram on the right. Thus, the definition of $\tilde{c}_{GG}$ in Eq. (35), with only $f=t$ explains the result of Eq. (20). Similarly, if the $c_{f}$ contributions are also considered, the relation Eq. (35) with $f\neq q$ explains Eq. (21) and the behaviour at high $p_{T}(a)$.
@@ 4.1 Summary of our findings
Overall, we see that precision measurements yield comparable or better bounds on the top-philic ALP than existing resonance searches. That said, such constraints are still rather mild and therefore mainly apply to the case of a relatively strongly coupled ALP. This shows how elusive this type of particle is and motivates further study into constraining top-philic ALPs, by either employing an even more global approach, or identifying new channels through which they can be constrained.
The interested reader can find much more detail in the next Secs. 4.24.5. We recall here that the case of a top-philic ALP decaying predominantly into invisible DM and therefore leading different phenomenology is investigated in Sec. 5.
@@ 4.4 Top-philic ALP in $t\bar{t}t\bar{t}$
[…]
@@
The contributions of each order in Eq. (41), excluding the SM one, are displayed in Fig. 10, for the total cross section at 13 TeV, setting $f_{a}/c_{t}=1\leavevmode\nobreak\ {\rm TeV}$. We explicitly denote them via their powers of $c_{t}$, $\alpha_{S}^{m}$ and $\alpha_{\rm EW}$. When the powers of the last two are not specified, the sum over the different $m$ and $l$ combinations for a given $n$ in Eq. (41) is understood. The contribution proportional to $c_{t}^{2}$ is clearly dominant, featuring a cancellation between a negative, purely QCD-mediated ($\alpha_{S}^{3}$) contribution and a positive, mixed QED-QCD mediated ($\alpha_{S}^{2}\alpha_{\rm EW}$) one. The contribution proportional to $c_{t}^{4}$ is entirely dominated by the purely QCD-mediated ($\alpha_{S}^{2}$) term. Although it may appear as very subleading w.r.t. the one proportional to $c_{t}^{2}$, we will show that, given our sensitivity to $f_{a}/c_{t}$, this contribution is also relevant.
@@ 4.5 Top-philic ALP in $t\bar{t}$
New states coupled to the top quark can also modify production rates via one-loop corrections, as has already been investigated in Refs. Esser:2023fdo ; Bruggisser:2023npd for the case of an ALP in top-quark pair production. In this section we present the first exact calculation of one-loop induced $c_{t}$ effects on the $t\bar{t}$ production cross section. We provide numerical results for various differential distributions and we use them to constrain the coupling of the top-philic ALP. Many technical details regarding this calculation are left for a future publication maltoni:2024 , where we will also provide analytical formulae.
@@
The quantity $\sigma_{\rm NP,\,virt}$ has previously been estimated by considering only the $s$-channel diagram in Fig. 11$(a)$ using an effective $aG\tilde{G}$ interaction dressed by an approximated form-factor Esser:2023fdo . In Ref. Bruggisser:2023npd the $s$-channel diagram has been calculated exactly, but the remaining loop diagrams from Fig. 11 have been only estimated. In our calculation all of the one-loop diagrams of order $(c_{t}/f_{a})^{2}$, as those in Fig. 11, are exactly taken into account. These diagrams in general contain UV divergencies, which have to be renormalised via the counterterms for the top-quark mass and its wave function. We perform renormalisation in the on-shell scheme, so no renormalisation scale dependence is present for what concerns $m_{t}$. We note in passing, that the higher-order term of Eq. (34) generates a one-loop correction to the self-energy of the top quark, see e.g. Ref. Galda:2023qjx . However, the corresponding diagram is a tadpole and therefore does not depend on the $p^{2}$ of the top quark. It amounts to a universal constant that can be reabsorbed into the definition of the top mass on-shell, bearing no physical consequences. More details on the full computation will be given in Ref. maltoni:2024 . In conclusion, since we exactly calculate all of the one-loop contributions, further effects w.r.t. Refs. Esser:2023fdo ; Bruggisser:2023npd are taken into account, such as new kinematic dependencies and also contributions from the $q\bar{q}$-initiated channels (Fig. 11$(c)$).
@@
We report representative results in Tab. 6 and, as already mentioned, the limits on $f_{a}/c_{t}$ are nearly independent of the ALP mass, which reflects how insensitive the virtual ALP corrections are to this quantity. In Fig. 8, the bounds from $t\bar{t}$ production correspond to the interpolation of the data reported in in Tab. 6. As discussed in Sec. 4.1, top-quark pair production data leads to the strongest constraints on $c_{t}$ in our model. Given the shape of the corrections induced by $\sigma_{\rm NP}/\sigma_{\rm SM}$, see Fig. 14, we expect that a finer binning close to the threshold would lead to stronger constraints. We encourage experimental collaborations to follow this strategy to further improve the bounds provided in this paper.
@@
[…]

### v2

@@ 1 Introduction
Axion-like Particles (ALPs) represent a compelling benchmark scenario for physics beyond the Standard Model (BSM) in which a new, singlet pseudoscalar state ($a$) is introduced in the low-energy spectrum of our theory of fundamental interactions. In particular, the ALP interactions with the Standard Model (SM) particles can be described within an effective field theory (EFT) that begins at canonical dimension-five, and respects an approximate shift symmetry, $a\to a+\textit{c}$ where $c$ is a constant, reflecting the fact that $a$ is the pseudo-Nambu Goldstone boson (pNGB) of a spontaneously broken global symmetry. As such, the ALP is naturally light compared to other BSM states, and can in fact be considered as the sole additional, dynamical degree of freedom at low energy.
Beyond the original motivation of the axion solution to the strong CP problem Peccei:1977hh; Peccei:1977ur; Wilczek:1977pj; PhysRevLett.48.11, ALPs can arise in many BSM scenarios such as string compactifications Cicoli:2013ana; Ringwald:2012cu, supersymmetric theories  Bellazzini:2017neg, neutrino-mass-generation mechanisms Alexandre:2020tba; Mavromatos:2020hfy, as well as composite-Higgs realisations featuring additional pNGBs Gripaios:2009pe. In addition, ALPs can be viable dark matter candidates as well as mediators to the dark-matter sector, see e.g. Ref. Preskill:1982cy; ABBOTT1983133; Dine:1982ah; Dror:2023fyd; Fitzpatrick:2023xks.
@@
Without committing to a specific UV completion, we will study the phenomenology of the top-philic ALP at the LHC, focusing on the model-independent features that primarily depend on the low-energy physics described by the EFT (see Ebadi:2019gij; Carmona:2021seb; Carmona:2022jid; Esser:2023fdo; Rygaard:2023dlx; Bruggisser:2023npd for recent related studies). The first observation is that the top-philic nature of the ALP is not preserved beyond tree level. In fact $c_{t}$ induces, through radiative corrections involving SM fields, couplings to all SM fermions and gauge bosons. As we will explore and discuss in detail, this has important implications for the possible discovery channels of this light BSM state. For example, for ALP masses below the bottom threshold, the constraints on the model are dominated by flavour observables (see, e.g., Ref. Bauer:2021mvw) from the induced fermion couplings. At even lighter, keV-scale masses, the induced coupling to electrons implies extremely strong bounds from stellar cooling via ALP emission, beyond any conceivable collider sensitivity Chala:2020wvs. Above the top-quark pair threshold, a top-philic ALP can be probed efficiently by resonant $t\bar{t}$ searches. However, by this point the ALP can no longer be considered light and shift-symmetry-breaking effects start to become significant.
@@ 2.1 The ALP effective Lagrangian
$\psi_{f}\rightarrow{\rm exp}\left(ic\frac{a(x)}{f_{a}}Q_{f}\right)\psi_{f},\quad\phi\rightarrow{\rm exp}\left(ic\frac{a(x)}{f_{a}}Q_{\phi}\right)\phi.$ (3)
@@ 2.2.2 Coupling to gluons and photons
The large suppression in Eq. (13) for relatively light ALPs raises the question of whether higher-loop contributions involving lighter fermions can actually become the dominant ones for $m_{a}\ll m_{t}$. While a full two-loop computation of $\mathcal{A}^{\mu\nu}(a\rightarrow gg)$ is beyond the scope of this work, we can estimate the size of this contribution. To this end, we consider the one-loop diagrams involving light fermions that couple to the ALP with the effective (one-loop) couplings in Eq. (9) and Eq. (10). These contributions correspond to diagrams such as those in Fig. 2 and Fig. 2. Thus, the effective interaction of Eq. (13) generalises to
@@
For the range we will consider in our work, $10\leavevmode\nobreak\ {\rm GeV}\lesssim m_{a}\lesssim 200\leavevmode\nobreak\ {\rm GeV}$, Eq. (17) shows that contributions from up, down and strange quarks are negligible, while the charm mass is relevant only in the proximity of $m_{a}\sim 10\leavevmode\nobreak\ {\rm GeV}$. Also, in the range $2m_{b}\ll m_{a}\ll 2m_{t}$, combining Eqs. (9), (16) and (17) one obtains
@@
First we notice, that at small $m_{a}$ values the bottom-quark contribution, with $c_{b}$ obtained from one-loop induced $c_{t}$ effects, is dominant w.r.t. the top-quark one. For larger $m_{a}$, the top contribution takes over. Second, in the mass-window of phenomenological interest in our paper, $10\leavevmode\nobreak\ {\rm GeV}\lesssim m_{a}\lesssim 200\leavevmode\nobreak\ {\rm GeV}$, $c_{GG}^{\rm eff}$ never surpasses 0.05, contrary to the $O(1)$ expectations from typical ALP models. Finally, there is a cancellation between the top and bottom contributions, arising at one- and two-loop level, respectively. We will discuss this in Sec. 3.1 and in Appendix E, noting that the exclusion limits that will be derived in this work are independent of a precise computation of this effect.
@@ 3.1 Branching ratios
The decay into photons follows a pattern very similar to the one discussed for the case of the decay into gluons. However, in this case, the loop-induced contributions involve electric charges and hence the minimum of the BR occurs at a numerically different value of $m_{a}$ w.r.t. the decay into gluons. The same argument regarding the location of the minimum and the irrelevance for our analysis applies to this decay channel. In particular, the relevant information for our analysis is that the BR of this channel is always (much) smaller than 0.1%. We also observe that, unlike the case of the gluons, there is no small increase in the region close to $m_{a}\approx 2m_{b}$; rather, there is an even stronger decrease, due to the additional dependence on the electric charge of the light fermion in the loop.
We now comment on the impact of the decay channels $a\rightarrow WW^{*}$ and $a\rightarrow ZZ^{*}$, with the subsequent decay of the gauge bosons into fermions. This case cannot be derived in general from Eq. (25), since in the derivation of those formulae both of the vector bosons were assumed to be on-shell; for $m_{a}<2m_{V}$, at least one $V$ has to be off-shell. Moreover, the ALPTopNLO_3flav UFO model presented in Appendix C cannot be used to perform these calculations, because the required ingredients for generating processes involving loops of EW gauge bosons have not been implemented. On the other hand, the impact of such decay modes can be estimated by comparing the case of the ALP with that of the SM Higgs boson, which has been studied in Ref. Denner:2011mq for various values of $m_{H}$. For the $ZZ^{*}$ case, the Higgs branching ratio for one virtual $Z^{*}$ ($m_{H}<2m_{Z}\simeq 180$ GeV) never surpasses the value of the branching ratio at $m_{H}=200\,\textrm{GeV}$, where both $Z$ bosons are on-shell. We expect a similar behaviour for the ALP, such that in the entire $m_{a}$ range considered the branching ratio will never exceed $0.01\%$, the value of BR($a\rightarrow ZZ$) for $m_{a}=200$ GeV. Furthermore, it is expected to decrease rather quickly for lower masses.
[…]
@@ 3.2 LHC production cross sections
Process Sub-processes $(Q=c,b)$ Scale Cuts
$pp\rightarrow a$ $gg\rightarrow a$, $Q\bar{Q}\rightarrow a$ $m_{a}$ None
$pp\rightarrow a+j$ $gg\rightarrow ga$, $gq\rightarrow qa$, $q\bar{q}\rightarrow ga$, $Q\bar{Q}\rightarrow ga$ $m_{\rm T}(a)$ $p_{\rm T}(a)>30\,{\rm GeV},\,|\eta(a)|<4$
$pp\rightarrow a+j_{h}$ $gQ\rightarrow Qa$ $m_{\rm T}(a)$ $p_{\rm T}(a)>30\,{\rm GeV},\,|\eta(a)|<2.5$
$pp\rightarrow t\bar{t}a$ $gg\rightarrow t\bar{t}a$, $q\bar{q}\rightarrow t\bar{t}a$ $m_{t}+\frac{m_{a}}{2}$ None
@@
As in the case of $gg\rightarrow a$, in addition to the contributions induced by $c_{t}$, the contributions from $c_{f}$ associated to $f\neq t$ cannot be neglected. These correspond to an approximate two-loop calculation that we obtain by plugging in the $c_{f}$ generated at one-loop. Concerning the change of basis from derivative to non-derivative, it is crucial to ensure that the definition of $\tilde{c}_{GG}$ in Eq. (35) includes $c_{f}$ for $f\neq t$ consistently with the $f$ running in the loop diagrams such as the one on the left of Fig. 5. Unlike the case of direct ALP production or its decay into gluons or photons, however, predictions for the $pp\rightarrow a+j$ production are much more reliable. Indeed, at high $p_{T}(a)$, which is the phenomenologically relevant kinematic configuration in our work, one expects a contribution from the top-quark loops of the form given in Eqs. (19) and (20) and a contribution from the other fermions of the form given in (21). Notice that, in the non-derivative basis, this means that the purely pseudoscalar component, associated to the left-hand diagram in Fig. 5 vanishes and all of the contribution is given by the tree-level diagram on the right. Thus, the definition of $\tilde{c}_{GG}$ in Eq. (35), with only $f=t$ explains the result of Eq. (20). Similarly, if the $c_{f}$ contributions are also considered, the relation Eq. (35) with $f\neq q$ explains Eq. (21) and the behaviour at high $p_{T}(a)$.
@@ 4.1 Summary of our findings
Overall, we see that precision measurements yield comparable or better bounds on the top-philic ALP than existing resonance searches. That said, such constraints are still rather mild and therefore mainly apply to the case of a relatively strongly coupled ALP. This shows how elusive this type of particle is and motivates further study into constraining top-philic ALPs, by either employing an even more global approach, or identifying new channels through which they can be constrained. The interested reader can find many more details in the next Secs. 4.24.5. We recall here that the case of a top-philic ALP decaying predominantly into invisible DM and therefore leading different phenomenology is investigated in Sec. 5.
Given the relatively small bound on $f_{a}/c_{t}$ from Eq. (38), one may wonder if the EFT approach is still valid, i.e., whether the energies probed are too high compared to the allowed scale of NP. This question has been addressed in detail in Esser:2023fdo and we briefly recap the main argument here. The validity of the EFT requires $f_{a}>Q(b)$ where $Q(b)$ is the typical scale of the process from which a bound $f_{a}/c_{t}<b$ is extracted. Whenever $b>Q(b)$ we are always in the range of validity of the EFT for couplings of order one or larger. Our bounds in fact correspond to the opposite region, in which $b\lesssim Q(b)$. However, the main point to keep in mind here is that we never bound $f_{a}$ alone but always its ratio with $c_{t}$. The validity of the EFT is then only preserved for relatively large values of $c_{t}\gtrsim 1$, which cannot realistically exceed $\mathcal{O}(4\pi)$. Considering the above, we can summarise the validity criterion by the following condition:
$\left\{\begin{split}f_{a}&\geq Q(b)\\ \frac{f_{a}}{c_{t}}&\geq b\\ \end{split}\right.\quad\longrightarrow\quad 4\pi>c_{t}\geq\frac{Q(b)}{b}\,,$ (39)
which is always satisfied for the analyses performed in this work.
@@ 4.4 Top-philic ALP in $t\bar{t}t\bar{t}$
$\mathcal{O}\left(\left(\frac{c_{t}}{f_{a}}\right)^{2n}\alpha_{S}^{m}\alpha_{\rm EW}^{l}\right)\qquad{\rm with}\leavevmode\nobreak\ n+m+l=4\leavevmode\nobreak\ {\rm and}\leavevmode\nobreak\ n\leq 2\,,$ (42)
@@
The contributions of each order in Eq. (42), excluding the SM one, are displayed in Fig. 10, for the total cross section at 13 TeV, setting $f_{a}/c_{t}=1\leavevmode\nobreak\ {\rm TeV}$. We explicitly denote them via their powers of $c_{t}$, $\alpha_{S}$ and $\alpha_{\rm EW}$. When the powers of the last two are not specified, the sum over the different $m$ and $l$ combinations for a given $n$ in Eq. (42) is understood. The contribution proportional to $c_{t}^{2}$ is clearly dominant, featuring a cancellation between a negative, purely QCD-mediated ($\alpha_{S}^{3}$) contribution and a positive, mixed QED-QCD mediated ($\alpha_{S}^{2}\alpha_{\rm EW}$) one. The contribution proportional to $c_{t}^{4}$ is entirely dominated by the purely QCD-mediated ($\alpha_{S}^{2}$) term. Although it may appear as very subleading w.r.t. the one proportional to $c_{t}^{2}$, we will show that, given our sensitivity to $f_{a}/c_{t}$, this contribution is also relevant.
@@ 4.5 Top-philic ALP in $t\bar{t}$
New states coupled to the top quark can also modify production rates via one-loop corrections, as has already been investigated in Refs. Esser:2023fdo; Bruggisser:2023npd for the case of an ALP in top-quark pair production. In this section we present the first exact calculation of one-loop induced $c_{t}$ effects on the $t\bar{t}$ production cross section. We provide numerical results for various differential distributions and we use them to constrain the coupling of the top-philic ALP. Many technical details regarding this calculation are left for a future publication Maltoni:2024wyh, where we will also provide analytical formulae.
@@
The quantity $\sigma_{\rm NP,\,virt}$ has previously been estimated by considering only the $s$-channel diagram in Fig. 11$(a)$ using an effective $aG\tilde{G}$ interaction dressed by an approximated form-factor Esser:2023fdo. In Ref. Bruggisser:2023npd the $s$-channel diagram has been calculated exactly, but the remaining loop diagrams from Fig. 11 have been only estimated. In our calculation all of the one-loop diagrams of order $(c_{t}/f_{a})^{2}$, as those in Fig. 11, are exactly taken into account. These diagrams in general contain UV divergencies, which have to be renormalised via the counterterms for the top-quark mass and its wave function. We perform renormalisation in the on-shell scheme, so no renormalisation scale dependence is present for what concerns $m_{t}$. We note in passing, that the higher-order term of Eq. (34) generates a one-loop correction to the self-energy of the top quark, see e.g. Ref. Galda:2023qjx. However, the corresponding diagram is a tadpole and therefore does not depend on the $p^{2}$ of the top quark. It amounts to a universal constant that can be reabsorbed into the definition of the top mass on-shell, bearing no physical consequences. More details on the full computation will be given in Ref. Maltoni:2024wyh. In conclusion, since we exactly calculate all of the one-loop contributions, further effects w.r.t. Refs. Esser:2023fdo; Bruggisser:2023npd are taken into account, such as new kinematic dependencies and also contributions from the $q\bar{q}$-initiated channels (Fig. 11$(c)$).
@@
We report representative results in Tab. 6 and, as already mentioned, the limits on $f_{a}/c_{t}$ are nearly independent of the ALP mass, which reflects how insensitive the virtual ALP corrections are to this quantity. In Fig. 8, the bounds from $t\bar{t}$ production correspond to the interpolation of the data reported in Tab. 6. As discussed in Sec. 4.1, top-quark pair production data leads to the strongest constraints on $c_{t}$ in our model. Given the shape of the corrections induced by $\sigma_{\rm NP}/\sigma_{\rm SM}$, see Fig. 14, we expect that a finer binning close to the threshold would lead to stronger constraints. We encourage experimental collaborations to follow this strategy to further improve the bounds provided in this paper.
@@
Figure 14: Differential $m(t\bar{t})$ (left) and $p_{\rm T}(t)$ (right) distributions similar to Fig. 13 but for different values of $m_{a}$ at $c_{t}/f_{a}=1$ TeV-1. Only the exact calculation is considered here unlike with Fig. 13.